OpenStudy (anonymous):

Product of (1+i)(1-2i)

OpenStudy (anonymous):

can you FOIL binomials?

OpenStudy (anonymous):

yes thats all i do?

OpenStudy (anonymous):

yes... that's it. then simplify.

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

ok that didnt work

OpenStudy (anonymous):

what did you get?

OpenStudy (lgbasallote):

remember that \(i^2 = -1\)

OpenStudy (anonymous):

2i^2-i+1

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

right...

OpenStudy (lgbasallote):

and it should be \(-2i^2 - i + 1\) coz \(-2i \times i = 2i^2\) sorry dp :p

OpenStudy (anonymous):

hahaha

OpenStudy (lgbasallote):

that's \(-2i \times i = -2i^2\)

OpenStudy (anonymous):

is the answer 1+4i?

OpenStudy (anonymous):

i really have no idea lol

OpenStudy (lgbasallote):

how did you get that? it's \(-2i^2 - i + 1\) like i said \(i^2 = -1\) so \((-2)(-1) - i +1 \)

OpenStudy (anonymous):

oh. lol i see

OpenStudy (anonymous):

i was just guessing

OpenStudy (anonymous):

i do that often trust me